OpenStudy (anonymous):

What is the sum of the arithmetic sequence 137, 125, 113 …, if there are 38 terms?

OpenStudy (anonymous):

here the common difference is 125-137 = -12 u use the formula Sn = n/2 [2a + (n-1)d] where n = 38, d = -12 and a = first term = 137

OpenStudy (anonymous):

are u happy with calculating that?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

ok